Is the measure of an angle is 2x-10, what is the measure of its complement?
Answer by stanbon(75887)

Is the measure of an angle is 2x-10, what is the measure of its complement?
-----
complement = 90-(2x-10) = 100-2x
=========================================
Cheers,
Stan H.
=========================================